OpenStudy (anonymous):

Please & thank you! Use properties of logarithms to solve the equation. log(base3) x + log(base3)(x + 2) = log(base3) 2 + log(base3) 12

OpenStudy (anonymous):

as i can recall log(ab)=loga+logb use this property

OpenStudy (anonymous):

i dont understand how to solve it though

mathslover (mathslover):

\[\large{log_3 x+log_3 (x+2)=log_3 2+log_3 {12}}\]

mathslover (mathslover):

\[\large{log_3(x(x+2))=log_3(2*12)}\]

mathslover (mathslover):

x(x+2) = 24

mathslover (mathslover):

can you solve further?

OpenStudy (anonymous):

would the answer just be be one?

mathslover (mathslover):

just tell me what is x(x+2)

OpenStudy (anonymous):

x^2 + 2x

mathslover (mathslover):

\[\large{x^2+2x-24=0}\] \[\large{x^2+6x-4x-24=0}\] \[\large{x(x+6)-4(x+6)=0}\] \[\large{(x-4)(x+6)=0}\]

mathslover (mathslover):

hence we have x = 4 or x = -6

OpenStudy (anonymous):

why is it -24 and not just 24

mathslover (mathslover):

x^2+2x=24 right?

mathslover (mathslover):

now subtract 24 both sides x^2+2x-24=24-24 x^2+2x-24=0

OpenStudy (anonymous):

oh okay! thank you!

mathslover (mathslover):

no problem and best of luck

OpenStudy (anonymous):

thanls

OpenStudy (anonymous):

but remember only the positive answer is accepted

OpenStudy (anonymous):

would it be wrong if i gave the negative answer as well?

mathslover (mathslover):

yes